# Understanding Video Transformers via Universal Concept Discovery

### v1

@@ Abstract
This paper studies the problem of concept-based interpretability of transformer representations for videos. Concretely, we seek to explain the decision-making process of video transformers based on high-level, spatiotemporal concepts that are automatically discovered. Prior research on concept-based interpretability has concentrated solely on image-level tasks. Comparatively, video models deal with the added temporal dimension, increasing complexity and posing challenges in identifying dynamic concepts over time. In this work, we systematically address these challenges by introducing the first Video Transformer Concept Discovery (VTCD) algorithm. To this end, we propose an efficient approach for unsupervised identification of units of video transformer representations - concepts, and ranking their importance to the output of a model. The resulting concepts are highly interpretable, revealing spatio-temporal reasoning mechanisms and object-centric representations in unstructured video models. Performing this analysis jointly over a diverse set of supervised and self-supervised representations, we discover that some of these mechanism are universal in video transformers. Finally, we demonstrate that VTCD can be used to improve model performance for fine-grained tasks.
@@ 1 Introduction
Figure 1: Heatmap predictions of the TCOW model [62] for tracking through occlusions (top), together with concepts discovered by our VTCD (bottom). We can see that the model encodes positional information in early layers, identifies containers and collisions events in mid-layers and tracks through occlusions in late layers. Only one video is shown, but the discovered concepts are shared between many dataset samples (see video for full results).
Understanding the hidden representations within neural networks is essential for addressing regulatory concerns [11, 31], preventing harms during deployment [5, 30], and can aid innovative model designs [13]. This problem has been studied extensively in the image world, both for convolutional neural networks (CNNs) [4, 35, 26, 22] and, more recently, for transformers [51, 64], resulting in a number of key insights. For example, image classification models extract low-level positional and texture cues at early layers and gradually combine them into higher-level, semantic concepts at later layers [4, 45, 24].
However, while video transformers do share their overall architecture with image-level ViTs, the insights obtained in existing works do very little to explain their inner mechanisms. Consider, for example, the recent approach for tracking occluded objects [62] shown in Figure 1 (top). To accurately reason about the trajectory of the invisible object inside the pot, texture or semantic cues alone would not suffice. What, then, are the spatiotemporal mechanisms used by this approach? And are any of these mechanisms universal across video models trained for different tasks?
To answer these questions, in this work we propose the Video Transformer Concept Discovery algorithm (VTCD) - the first concept-discovery methodology for interpreting the representations of deep video transformers. We focus on concept-based interpretability [26, 69, 22, 21] due to its capacity to explain the decision-making process of a complex model’s distributed representations in high-level, intuitive terms. Our goal is to decompose a representation at any given layer into human-interpretable ‘concepts’ without any labelled data (i.e., concept discovery) and then rank them in terms of their importance to the model output.
Concretely, we first group model features at a given layer into spatio-temporal tubelets via SLIC clustering [1], which serve as a basis for our analysis (Section 3.1.1). Next, we cluster these tubelets across videos to discover high-level concepts [14, 39, 69, 22, 21] (Section 3.1.2) . The resulting concepts for an occluded object tracking method [62] are shown in Figure 1 (bottom) and span a broad range of cues, including spatiotamporal ones that detect events, like collisions, or track the containers.
@@
Next, we use VTCD to study whether there are any universal mechanisms in video transformer models, that emerge irrespective of their training objective. To this end, we extend the recent work by Dravid et al. [16] to automatically identify important concepts that are shared between several models in Section 4.1. We then analyze a diverse set of representations (e.g. supervised, self-supervised, or video-language) and make a number of discoveries: (i) many concepts are indeed shared between models trained for different tasks; (ii) early layers tend to form a spatiotemporal basis that underlines the rest of the information processing; (iii) later layers form object-centric video representations even in models trained in a self-supervised way.
Finally, VTCD can be used to turn a pre-trained video transformer into an efficient and effective fine-grained recognition model by pruning least important units. We demonstrate in Section 5.4 that removing one third of the heads from an action classification model results in a 4.3% increase in accuracy while reducing computation by 33%.
@@ 2 Related work
Our work proposes a novel concept-based interpretability algorithm that focuses on transformer-based representations for video understanding. Below, we review the most relevant works in each of these fields.
@@
In contrast, unsupervised concept discovery makes no assumptions on the existence of semantic concepts and uses clustering to partition data into interpretable components within the model’s feature space. ACE [26] and CRAFT [22] segment input images into superpixels and random crops, before applying clustering at a given layer. In videos, however, the potential tubelets far outnumber image crops, prompting us to introduce a more efficient method for segmenting videos into proposals in Section 3.1.1.
@@
Recent work [16] identifies neurons that produce similar activation maps across various image models (including transformers). However, neurons are unable to explain a full extent of a models’ representation due to it being distributed across many dimensions [18]. In contrast, our method works on any dimensional features and is applied to video-based models.
Interpretability of transformers has received a significant amount of attention recently, due to the success of this architecture in variety of computer vision problems. Early work [51] contrasted vision transformer representations with CNNs (representational differences per layer, receptive fields, localization of information, etc). Other work aims to generate saliency heatmaps based on attention maps of a model [9, 8]. Later works focused on understanding the impact of different training protocols [47, 64] (e.g. self-supervised learning (SSL) vs. supervised) and robustness [70, 50]. The features of a specific SSL vision transformer, DINO [2], were explored in detail and shown to have surprising utility for part-based segmentation tasks. However, none of these works address concept-based interpretability or study video representations.
Independently, studies in natural language processing (NLP) have analyzed self-attention layers [17, 63] and found that heads are often specialized to capture different linguistic or grammatical phenomenon. This is qualitatively seen in works that shows dissimilar attention maps for different self-attention heads [12, 34]. Moreover, other NLP works [44, 63] explore the impact of removing heads and find that only a small number need to be kept to produce similar performance. Our findings agree with evidence from these works, and in Section 5.4 we further demonstrate that pruning unimportant heads from video transformers can actually improve model’s performance.
@@ 3 Video transformer concept discovery
We study the problem of decomposing a video representation into a set of high-level open-world concepts and ranking their importance for the model’s predictions. We are given a set of (RGB) videos, $\textbf{X}\in\mathbb{R}^{N\times 3\times T\times H\times W}$, where $N$, $T$, $H$, and $W$ denote the dataset size, time, height, and width, respectively, and an $L$ layer pretrained model, $f$. Let $f_{[r,l]}$ denote the model from layer $r$ to $l$, with $f_{[1,l]}(\textbf{X})=\textbf{Z}_{l}\in\mathbb{R}^{N\times C\times T^{\prime}\times H^{\prime}\times W^{\prime}}$ being the intermediate representation at layer $l$. To decompose $\textbf{Z}_{l}$ into a set of human-interpretable concepts, $\textbf{C}_{l}=\{c_{1},...,c_{Q}\}$, existing, image-level approaches [26, 22] first parse the $N$ feature maps into a set of $M$ proposals, $\textbf{T}\in\mathbb{R}^{M\times C}$ ($M>N$), where each $T_{m}$ corresponds to a region of the input image. These proposals are then clustered into $Q<<M$ concepts in the feature space of the model to form an assignment matrix $W\in\mathbb{R}^{M\times Q}$. Finally, the importance of each concept $c_{i}$ to the model’s prediction is quantified by a score $s_{i}\in[0,1]$. Performing this analysis over all the layers in $f$ produces the entire set of concepts for a model, $\textbf{C}=\{\textbf{C}_{1},...,\textbf{C}_{L}\}$, together with their corresponding importance scores.
However, existing approaches are not immediately applicable to video transformers because they do not scale well and are focused on 2D CNN architectures. In this work, we extend concept-based interpretability to video representations. To this end, we first describe a computationally tractable proposal generation method (Section 3.1.1) that operates over space-time feature volumes and outputs spatiotemporal tublets. Next (Section 3.1.2), we adapt existing concept clustering techniques to video transformer representations. Finally, in Section 3.2 we propose CRIS - a novel concept importance estimation approach applicable to any architecture units, including transformer heads.
@@ 3.1.1 Tubelet proposals
Previous methods [26, 22, 21] use superpixels or crops in RGB space to propose segments; however, the number of possible segments is exponentially greater for videos. Moreover, proposals in color space are unrestricted and may not align with the model’s encoded information, leading to many irrelevant or noisy segments. To address these drawbacks, we instantiate proposals in feature space, which naturally partitions a video based on the information contained within each layer (shown in Figure 2, left).
More specifically, we construct tubelets per video via Simple Linear Iterative Clustering [1, 32] (SLIC) on the spatiotemporal features via
@@
where $\textbf{T}\in\mathbb{R}^{M\times C}$ is the set of tubelets for the dataset, $\textbf{B}\in\{0,1\}^{C\times M\times N\times T^{\prime}\times H^{\prime}\times W^{\prime}}$ are spatiotemporal binary support masks obtained from the SLIC clustering, $M$ is the total number of tubelets for all $N$ videos ($M>>N$), and GAP is a global average pooling operation over the space and time dimensions.
SLIC is an extension of the K-Means algorithm that controls a trade-off between cluster support regularity and adaptability, and also constrains cluster support masks to be connected. Together these properties produce non-disjoint tubelets that are easier to interpret for humans because they reduce the need to attend to multiple regions in a video at a time. Further, the pruning step in SLIC makes it more robust to the hyperparameter that controls the desired number of clusters, as it automatically prunes spurious, disconnected tubelets. Next, we describe our approach for grouping individual tubelets into higher-level concept clusters.
@@ 3.1.2 Concept clustering
where $\textbf{G}\in[0,1]^{C\times Q}$ and $\sum_{j}\textbf{G}_{i,j}=1$. To cluster a set of tublets, T, into corresponding concepts, we optimize
@@
where the final set of concepts are represented by the rows of the matrix C, i.e. concept centroid $c_{i}$ is the $i^{th}$ row of C (Figure 2, right).
@@ 3.2 Concept importance
However, while this approach works well for CNNs [21], transformers are robust to small perturbations within self-attention layers [44, 63]. Therefore, single concept masking has little effect on performance (shown by results in Figure 4). Instead, we mask a high percentage of sampled concepts in parallel (across all layers and heads) and then empirically validate in Section 5.1 that averaging the results over thousands of samples produces valid concept rankings.
Formally, we propose Concept Randomized Importance Sampling (CRIS), a robust method to compute importance for any unit of interest. To this end, we first randomly sample $K$ different concept sets, such that each $^{k}\subset\textbf{C}$ C. We then define $\textbf{C}^{k}_{l}$ as the set of concepts in $\textbf{C}^{k}$ discovered at layer $l$, with $\textbf{B}_{\textbf{C}^{k}_{l}}$ denoting the corresponding binary support masks. We mask out every concept at every layer of the model via
@@ 4 Understanding transformers with VTCD
Our algorithm facilitates the identification of concepts within any unit of a model and quantifying their significance in the final predictions. However, this is not enough to fully represent the computations performed by a video transformer. It is also crucial to understand how these concepts are employed in the model’s information flow.
As several recent works have shown [17, 46], the residual stream of a transformer serves as the backbone of the information flow. Each self-attention block then reads information from the residual stream with a linear projection, performs self-attention operations to process it, and finally writes the results back into the residual stream. Crucially, self-attention processing is performed individually for each head and several studies have shown, both in vision [2, 15, 34] and NLP [63], that different self-attention heads capture distinct information. In other word, heads form the basis of the transformer representation.
A closer analysis of the concepts found in the heads of TCOW with VTCD allows us to identify several patterns in the information processing of that model. In particular, Figure 1 shows that the heads in early layers group input tokens based on their spatiotemporal positions. This information is then used to track objects and identify events in mid-layers, and later layers utilize mid-layer representations to reason about occlusions. Next, we study whether any of these mechanisms are universal across video transformers trained on different datasets with varying objectives.
@@ 4.1 Rosetta concepts
Inspired by [16], we propose to mine for Rosetta concepts that are shared between models and represent the same information. The key to identifying Rosetta units is a robust metric, $R$, where a higher R-score corresponds to the two units having a larger amount of shared information. Previous work [16] focused on finding such neurons in image models based on correlating their activation maps. We instead propose to measure the similarity between concepts (i.e. distributed representations) via the mean Intersection over Union (mIoU) of the concepts’ support.
Formally, we mine Rosetta concepts by first applying VTCD to a set of $D$ models $\{f^{1},...,f^{D}\}$, resulting in discovered concepts, $\textbf{C}^{j}=\{c^{j}_{1},...,c^{j}_{i}\}$, and importance scores, $\textbf{S}^{j}=\{s^{j}_{1},...,s^{j}_{i}\}$, for each model $f^{j}$. We then aim to measure the similarity between all concept $D$-tuples from the models. Given a set of $D$ concepts, $\{c^{1}_{i},...,c^{D}_{i}\}$ and corresponding binary support masks, $\{\textbf{B}^{1}_{i},...,\textbf{B}^{D}_{i}\}$, we define the similarity score of these concepts as
@@
Naively computing the similarity between all $D$-tuples results in an exponential number of computations and is intractable for even small $D$. To mitigate this issues, we exclude two types of concepts: (i) unimportant ones and (ii) those with a low R-score among $d$-tuples where $d<D$. More specifically, we only consider the most important $\epsilon\%$ of concepts from each model. We then iterate over $d\in\{2,...,D\}$ and filter out any concept that has an R-score less than $\delta$ for all d-tuples in which it participates. Formally, the filtered Rosetta d-concept scores are defined as
@@
where $\textbf{R}^{d}_{\epsilon}$ is the set of all R-scores among $d$ concepts after the $\epsilon$ importance filtering. This results in a significantly smaller pool of candidates for the next stage $d+1$, reducing the overall computational complexity of the algorithm. Finally, as some concepts may reside in a subset of the models but are still interesting to study, we examine the union of all important and confident Rosetta d-concepts corresponding to R-scores $\textbf{R}^{2}_{\epsilon,\delta}\cup\cdot\cdot\cdot\cup\textbf{R}^{D}_{\epsilon,\delta}$.

### v2

@@ Abstract
This paper studies the problem of concept-based interpretability of transformer representations for videos. Concretely, we seek to explain the decision-making process of video transformers based on high-level, spatiotemporal concepts that are automatically discovered. Prior research on concept-based interpretability has concentrated solely on image-level tasks. Comparatively, video models deal with the added temporal dimension, increasing complexity and posing challenges in identifying dynamic concepts over time. In this work, we systematically address these challenges by introducing the first Video Transformer Concept Discovery (VTCD) algorithm. To this end, we propose an efficient approach for unsupervised identification of units of video transformer representations - concepts, and ranking their importance to the output of a model. The resulting concepts are highly interpretable, revealing spatio-temporal reasoning mechanisms and object-centric representations in unstructured video models. Performing this analysis jointly over a diverse set of supervised and self-supervised representations, we discover that some of these mechanism are universal in video transformers. Finally, we show that VTCD can be used for fine-grained action recognition and video object segmentation.
@@ 1 Introduction
Figure 1: Heatmap predictions of the TCOW model [66] for tracking through occlusions (top), together with concepts discovered by our VTCD (bottom). We can see that the model encodes positional information in early layers, identifies containers and collision events in mid-layers and tracks through occlusions in late layers. Only one video is shown, but the discovered concepts are shared between many dataset samples (see video for full results).
Understanding the hidden representations within neural networks is essential for addressing regulatory concerns [14, 34], preventing harms in deployment [8, 33], and can aid innovative model designs [16]. This problem has been studied extensively for images, both for convolutional neural networks (CNNs) [5, 37, 29, 25] and, more recently, vision transformers (ViTs) [55, 68], resulting in multiple key insights. For example, image classification models extract low-level positional and texture cues at early layers and gradually combine them into higher-level, semantic concepts at later layers [5, 49, 27].
However, while video transformers do share their overall architecture with image-level ViTs, the insights obtained in existing works do very little to explain their inner mechanisms. Consider, for example, the recent approach for tracking occluded objects [66] shown in Figure 1 (top). To accurately reason about the trajectory of the invisible object inside the pot, texture or semantic cues alone would not suffice. What, then, are the spatiotemporal mechanisms used by this approach? Are any of these mechanisms universal across video models trained for different tasks?
To answer these questions, in this work we present the Video Transformer Concept Discovery algorithm (VTCD) - the first concept discovery methodology for interpreting the representations of deep video transformers. We focus on concept-based interpretability [29, 74, 25, 24] due to its capacity to explain the decision-making process of a complex model’s distributed representations in high-level, intuitive terms. Our goal is to decompose a representation at any given layer into human-interpretable ‘concepts’ without any labelled data (i.e. concept discovery) and then rank them in terms of their importance to the model output.
Concretely, we first group model features at a given layer into spatiotemporal tubelets via SLIC clustering [1], which serve as a basis for our analysis (Section 3.1.1). Next, we cluster these tubelets across videos to discover high-level concepts [17, 42, 74, 25, 24] (Section 3.1.2). The resulting concepts for an occluded object tracking method [66] are shown in Figure 1 (bottom) and span a broad range of cues, including spatiotemporal ones that detect events, like collisions, or track the containers.
@@
Next, we use VTCD to study whether there are any universal mechanisms in video transformer models, that emerge irrespective of their training objective. To this end, we extend recent work [19] to automatically identify important concepts that are shared between several models in Section 4.1. We then analyze a diverse set of representations (e.g. supervised, self-supervised, or video-language) and make a number of discoveries: (i) many concepts are indeed shared between models trained for different tasks; (ii) early layers tend to form a spatiotemporal basis that underlines the rest of the information processing; (iii) later layers form object-centric video representations, even in models trained in a self-supervised way.
We also show how VTCD can be applied for downstream tasks. Firstly, pruning the heads of an action classification model according to their estimated importance yields a 4.3% increase in accuracy while reducing computation by 33%. Secondly, object-centric concepts discovered by VTCD can be used for video-object segmentation (VOS) and achieve strong performance on the DAVIS’16 benchmark [52] even for self-supervised representations.
@@ 2 Related work
Our work introduces a novel concept-based interpretability algorithm that focuses on transformer-based representations for video understanding. Below, we review the most relevant works in each of these fields.
@@
In contrast, unsupervised concept discovery makes no assumptions on the existence of semantic concepts and uses clustering to partition data into interpretable components within the model’s feature space. ACE [29] and CRAFT [25] segment input images into superpixels and random crops, respectively, before applying clustering at a given layer. In videos, however, the potential tubelets far outnumber image crops, prompting us to introduce a more efficient method for segmenting videos into proposals in Section 3.1.1.
@@
Recent work [19] identifies neurons that produce similar activation maps across various image models (including transformers). However, neurons are unable to explain the full extent of a models’ distributed [21] representation. In contrast, our method works on features of arbitrary dimension and is applied to video models.
Interpretability of transformers has received significant attention recently, due to its success in a variety of computer vision tasks. Early work [55] contrasted vision transformer representations with CNNs (representational differences per layer, receptive fields, localization of information, etc). Other work aims to generate saliency heatmaps based on attention maps of a model [12, 11]. Later works focused on understanding the impact of different training protocols [51, 68] (e.g. self-supervised learning (SSL) vs. supervised) and robustness [75, 54]. The features of a specific SSL vision transformer, DINO [2], were explored in detail and shown to have surprising utility for part-based segmentation tasks. However, none of these works address concept-based interpretability or study video representations.
Independently, studies in natural language processing (NLP) analyzed self-attention layers [20, 67] and found that heads are often specialized to capture different linguistic or grammatical phenomenon. This is qualitatively seen in vision works that show dissimilar attention maps for different self-attention heads [15, 36]. Moreover, other NLP works [48, 67] explore the impact of removing heads and find that only a small number need to be kept to produce similar performance. Our findings agree with evidence from these works and in Section 5 we further demonstrate that targeted pruning of unimportant heads from video transformers can actually improve a model’s performance.
@@ 3 Video transformer concept discovery
We study the problem of decomposing a video representation into a set of high-level open-world concepts and ranking their importance for the model’s predictions. We are given a set of (RGB) videos, $\textbf{X}\in\mathbb{R}^{N\times 3\times T\times H\times W}$, where $N$, $T$, $H$, and $W$ denote the dataset size, time, height, and width, respectively, and an $L$ layer pretrained model, $f$. Let $f_{[r,l]}$ denote the model from layer $r$ to $l$, with $f_{[1,l]}(\textbf{X})=\textbf{Z}_{l}\in\mathbb{R}^{N\times C\times T^{\prime}\times H^{\prime}\times W^{\prime}}$ being the intermediate representation at layer $l$. To decompose $\textbf{Z}_{l}$ into a set of human-interpretable concepts, $\textbf{C}_{l}=\{c_{1},\ldots,c_{Q}\}$, existing, image-level approaches [29, 25] first parse the $N$ feature maps into a set of $M$ proposals, $\textbf{T}\in\mathbb{R}^{M\times C}$ ($M>N$), where each $T_{m}$ corresponds to a region of the input image. These proposals are then clustered into $Q<<M$ concepts in the feature space of the model to form an assignment matrix $W\in\mathbb{R}^{M\times Q}$. The importance of each concept $c_{i}$ to the model’s prediction is quantified by a score $s_{i}\in[0,1]$. Performing this analysis over all layers in $f$ produces the entire set of concepts for a model, $\textbf{C}=\{\textbf{C}_{1},\ldots,\textbf{C}_{L}\}$, together with their corresponding importance scores.
Existing approaches are not immediately applicable to video transformers because they do not scale well and are focused on 2D CNN architectures. In this work, we extend concept-based interpretability to video representations. To this end, we first describe a computationally tractable proposal generation method (Section 3.1.1) that operates over space-time feature volumes and outputs spatiotemporal tubelets. Next (Section 3.1.2), we adapt existing concept clustering techniques to video transformer representations. Finally, in Section 3.2 we propose a novel concept importance estimation approach applicable to any architecture units, including transformer heads.
@@ 3.1.1 Tubelet proposals
Previous proposal methods [29, 25, 24] use superpixels or crops in RGB space to generate segments; however, the number of possible segments is exponentially greater for videos. Moreover, proposals in color space are unrestricted and may not align with the model’s encoded information, leading to many irrelevant or noisy segments. To address these drawbacks, we instantiate proposals in feature space, which naturally partitions a video based on the information contained within each layer (shown in Figure 2, left).
We produce tubelets per-video via Simple Linear Iterative Clustering [1] (SLIC) on the spatiotemporal features as
@@
where $\textbf{T}\in\mathbb{R}^{M\times C}$ is the set of tubelets for the dataset, $\textbf{B}\in\{0,1\}^{C\times M\times N\times T^{\prime}\times H^{\prime}\times W^{\prime}}$ are spatiotemporal binary support masks obtained from SLIC, $M$ is the number of tubelets for all $N$ videos ($M>>N$), and GAP represents global average pooling over the space and time dimensions.
SLIC is an extension of the K-Means algorithm [46] that controls a trade-off between cluster support regularity and adaptability, and also constrains cluster support masks to be connected. Together these properties produce non-disjoint tubelets that are easier to interpret for humans because they reduce the need to attend to multiple regions in a video at a time. Further, the pruning step in SLIC makes it more robust to the hyperparameter that controls the desired number of clusters, as it automatically prunes spurious, disconnected tubelets. Next, we describe our approach for grouping individual tubelets into higher-level concept clusters.
@@ 3.1.2 Concept clustering
where $\textbf{G}\in[0,1]^{C\times Q}$ and $\sum_{j}\textbf{G}_{i,j}=1$. To cluster a set of tubelets, T, into corresponding concepts, we optimize
@@
where the final set of concepts are the rows of the matrix C, i.e. concept centroid $c_{i}$ is the $i^{th}$ row of C (Figure 2, right).
@@ 3.2 Concept importance
While this approach works well for CNNs [24], transformers are robust to small perturbations within self-attention layers [48, 67]. Therefore, single concept masking has little effect on performance (shown by results in Figure 4). Instead, we mask a high percentage of sampled concepts in parallel (across all layers and heads) and then empirically validate in Section 5.1 that averaging the results over thousands of samples produces valid concept rankings.
Formally, we propose Concept Randomized Importance Sampling (CRIS), a robust method to compute importance for any unit of interest. To this end, we first randomly sample $K$ different concept sets, such that each $^{k}\subset\textbf{C}$ C. We then define $\textbf{C}^{k}_{l}$ as the set of concepts in $\textbf{C}^{k}$ discovered at layer $l$, with $\textbf{B}_{\textbf{C}^{k}_{l}}$ denoting the corresponding binary support masks. We mask out every concept at every layer via
@@ 4 Understanding transformers with VTCD
Our algorithm facilitates the identification of concepts within any unit of a model and quantifying their significance in the final predictions; however, this is insufficient to fully represent the computations performed by a video transformer. It is also crucial to understand how these concepts are employed in the model’s information flow.
As several recent works have shown [20, 50], the residual stream of a transformer serves as the backbone of the information flow. Each self-attention block then reads information from the residual stream with a linear projection, performs self-attention operations to process it, and finally writes the results back into the residual stream. Crucially, self-attention processing is performed individually for each head with several studies showing, both in vision [2, 18, 36] and NLP [67], that different self-attention heads capture distinct information. In other word, heads form the basis of the transformer representation.
A closer analysis of the concepts found in the heads of the TCOW model [66] with VTCD allows us to identify several information processing patterns in that model. In particular, Figure 1 shows that the heads in early layers group input tokens based on their spatiotemporal positions. This information is then used to track objects and identify events in mid-layers, and later layers utilize mid-layer representations to reason about occlusions. Next, we study which of these mechanisms are universal across video transformers trained on different datasets and objectives (cf. [19]).
@@ 4.1 Rosetta concepts
Inspired by previous work [19], we mine for Rosetta concepts that are shared between models and represent the same information. The key to identifying Rosetta units is a robust metric, $R$, where a higher $R$-score corresponds to the two units having a larger amount of shared information. This previous work [19] focused on finding such neurons in image models based on correlating their activation maps. We instead measure the similarity between concepts (i.e. distributed representations) via the mean Intersection over Union (mIoU) of the concepts’ support.
Formally, we mine Rosetta concepts by first applying VTCD to a set of $D$ models $\{f^{1},\ldots,f^{D}\}$, resulting in discovered concepts, $\textbf{C}^{j}=\{c^{j}_{1},\ldots,c^{j}_{i}\}$, and importance scores, $\textbf{S}^{j}=\{s^{j}_{1},\ldots,s^{j}_{i}\}$, for each model $f^{j}$. We then aim to measure the similarity between all concept $D$-tuples from the models. Given a set of $D$ concepts, $\{c^{1}_{i},\ldots,c^{D}_{i}\}$, and corresponding binary support masks, $\{\textbf{B}^{1}_{i},\ldots,\textbf{B}^{D}_{i}\}$, we define the similarity score of these concepts as
@@
Naively computing the similarity between all $D$-tuples results in an exponential number of computations and is intractable for even small $D$. To mitigate these issues, we exclude two types of concepts: (i) unimportant ones and (ii) those with a low $R$-score among $d$-tuples, where $d<D$. More specifically, we only consider the most important $\epsilon\%$ of concepts from each model. We then iterate over $d\in\{2,\ldots,D\}$ and filter out concepts that have $R$-scores less than $\delta$ for all d-tuples in which it participates. Formally, the filtered Rosetta d-concept scores are defined as
@@
where $\textbf{R}^{d}_{\epsilon}$ is the set of all $R$-scores among $d$ concepts after the $\epsilon$ importance filtering. This results in a significantly smaller pool of candidates for the next stage, $d+1$, reducing the overall computational complexity of the algorithm. Finally, as some concepts may reside in a subset of the models but are still interesting to study, we examine the union of all important and confident Rosetta d-concepts corresponding to $R$-scores $\textbf{R}^{2}_{\epsilon,\delta}\cup\cdot\cdot\cdot\cup\textbf{R}^{D}_{\epsilon,\delta}$.